OpenStudy (anonymous):

How do determine that these vectors are not parallel? <-8, √3> and <-2√3, -16>

OpenStudy (anonymous):

I know they are orthogonal, but I dont know how to rule out parallel

OpenStudy (amistre64):

whats their dot product?

OpenStudy (anonymous):

0

OpenStudy (amistre64):

then they are perpendicular, and not parallel

OpenStudy (anonymous):

yeah but what is the operation i need to do if they are parallel, my book is convoluted

OpenStudy (amistre64):

i just do a ratio test

OpenStudy (amistre64):

<a,b,c> ------ <p,q,r> if the ratios equal then they are scalars a/p = b/q = c/r is parallel, or same line

OpenStudy (anonymous):

ok so in this instance -8/-2√3 must = √3/-16 in order to parallel?

OpenStudy (amistre64):

yes, or at least in order to have the chance of being parallel

OpenStudy (amistre64):

parallel vectors have scalar position vectors

OpenStudy (amistre64):

let u = <a,b,c> let v = nu = <na,nb,nc> a/na = 1/n b/nb = 1/n c/nc = 1/n

OpenStudy (anonymous):

Ok, I think I follow the logic. Thanks a lot.

OpenStudy (amistre64):

youre welcome